# Flexible Deep Neural Networks for Partially Linear Survival Data

### v1

@@ 8 Appendix - Proofs
Let $\mbox{${\theta}$}\in\mathbb{R}^{p}$ and let $\{g_{s}:s\in(-1,1)\}\subset\mathcal{H}_{g_{o}}$ be a one–dimensional regular submodel of $g$, with
$g_{s}=g_{o}+sg,\qquad g=\left.\frac{\partial g_{s}}{\partial s}\right|_{s=0}\in\mathcal{T}_{g_{o}}\,.$
@@
$\left\{P_{\theta_{o}+s\theta,\;g_{s}}:s\in(-1,1)\right\}\,.$
For a single observation $(T,\Delta,{\bf X},{\bf Z})$ the log-likelihood is
@@
Differentiating $\ell(\mbox{${\theta}$}_{o}+s\mbox{${\theta}$},g_{s})$ at $s=0$ gives
$\left.\frac{d}{ds}\ell(\mbox{${\theta}$}_{o}+s\mbox{${\theta}$},g_{s})\right|_{s=0}=\mbox{${\theta}$}^{\top}\dot{\ell}_{\theta_{o}}+\dot{\ell}_{g_{o},g},$
@@
$\dot{\ell}_{\theta_{o}}=\Delta{\bf Z}-\int_{0}^{\tau}{\bf Z}Y(t)\exp\{g_{o}(t,{\bf X})+\mbox{${\theta}$}_{o}^{\top}{\bf Z}\}\,dt=\int_{0}^{\tau}{\bf Z}\,dM(t),$
and
$\dot{\ell}_{g_{o},g}=\Delta g(T,{\bf X})-\int_{0}^{\tau}g(t,{\bf X})\,Y(t)\exp\{g_{o}(t,{\bf X})+\mbox{${\theta}$}_{o}^{\top}{\bf Z}\}\,dt=\int_{0}^{\tau}g(t,{\bf X})\,dM(t)\,.$
Thus the (naive) score for $\mbox{${\theta}$}_{o}$ is $\dot{\ell}_{\theta_{o}}$. To obtain the efficient score, we project this onto the orthocomplement of the nuisance tangent space
$\mathcal{T}_{g_{o}}^{\mathrm{nuis}}=\left\{\dot{\ell}_{g_{o},g}:g\in\mathcal{T}_{g_{o}}\right\},$
but the projection must be taken onto the closure $\overline{\mathcal{T}_{g_{o}}^{\mathrm{nuis}}}$ in $L^{2}(P_{0})$. Hence the efficient score is
@@
where $\Pi_{g_{o}}$ denotes the $L^{2}(P_{0})$–projection onto $\overline{\mathcal{T}_{g_{o}}^{\mathrm{nuis}}}$. The projection $\Pi_{g_{o}}(\dot{\ell}_{\theta_{o}})$ has the representation $\dot{\ell}_{g_{o},{\bf g}^{*}}$ for a pp–vector ${\bf g}^{*}=(g_{1}^{*},\dots,g_{p}^{*})^{\top}\in(\overline{\mathcal{T}_{g_{o}}})^{p}$ satisfying the orthogonality condition
$\mathbb{E}\!\Big\{(\dot{\ell}_{\theta_{o}}-\dot{\ell}_{g_{o},{\bf g}^{*}})\,\dot{\ell}_{g_{o},g}\Big\}={\bf 0}\qquad\text{for all }g\in\mathcal{T}_{g_{o}}\,.$
Using Lemma 1 of sasieni1992information, this orthogonality is equivalent to
$\mathbb{E}\!\left[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}\,g(T,{\bf X})\right]={\bf 0}\qquad\text{for all }g\in\mathcal{T}_{g_{o}}\,.$
Thus, componentwise, $g_{j}^{*}$ is the (unique) $\Delta$–weighted $L^{2}(P_{0})$ projection of $Z_{j}$ onto $\overline{\mathcal{T}_{g_{o}}}$. Equivalently, $\bf g$$^{*}$ minimizes
$\mathbb{E}\!\left(\Delta\|{\bf Z}-{\bf g}(T,{\bf X})\|_{c}^{2}\right)\,,\qquad{\bf g}\in(\overline{\mathcal{T}_{g_{o}}})^{p}\,.$
Substituting $\bf g$$^{*}$, the efficient score becomes
@@
By the martingale isometry (again Lemma 1 of sasieni1992information), the efficient information matrix is
$I(\theta_{o})=\mathbb{E}\left[\Delta\,\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}^{\otimes 2}\right].$
Finally, under the regularity conditions in Assumptions (A1)–(A4), existence and uniqueness of the weighted projection $\bf g$$^{*}$ in $\overline{\mathcal{T}_{g_{o}}}^{p}$ follow from the approximation theorems of stone1985additive and the general semiparametric projection theory in Appendix A.4 of bickel1993efficient. Therefore, both the efficient score and the efficient information matrix are well-defined and given by the formulas above. ∎
@@
where $\bf g$$^{*}$ is the pp–dimensional minimizer from Theorem 3, which depends only on $(\mbox{${\theta}$}_{o},g_{o})$. By Theorem 3, at the true parameter $(\mbox{${\theta}$}_{o},g_{o})$ this coincides with the efficient score,
@@
The empirical efficient score at $(\mbox{${\theta}$},g)$ is $\mathbb{P}_{n}r_{\theta,g}$. By the first–order condition for the (sieve) maximum–likelihood estimator, the estimated parameters satisfy
$\mathbb{P}_{n}r_{\widehat{\theta},\widehat{g}}={\bf 0}\,.$
Write
@@
$\mathbb{P}\bigl\{r_{\widehat{\theta},\widehat{g}}-r_{\theta_{o},g_{o}}\bigr\}=-\,P_{n}r_{\theta_{o},g_{o}}+o_{p}(n^{-1/2}).$ (23)
@@
Fix a deterministic direction $g$ and consider $g_{s}=g_{o}+sg$. Differentiating under the integral sign and applying the martingale compensation identity yields
$\left.\frac{d}{ds}\right|_{s=0}m(\mbox{${\theta}$}_{o},g_{s})=-\,\mathbb{E}\bigl[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}g(T,{\bf X})\bigr]\,.$
By Theorem 3, $\bf g$$^{*}$ is the $\Delta$–weighted $L^{2}(P)$–projection of $\bf Z$} onto the nuisance tangent space $(\overline{\mathcal{T}}_{g_{o}})^{p}$, so
$\mathbb{E}\bigl[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}g(T,{\bf X})\bigr]={\bf 0}\quad\text{for all }h\in\overline{\mathcal{T}}_{g_{o}}.$
By Assumptions A1A4, $\widehat{g}-g_{o}\in\overline{\mathcal{T}}_{g_{o}}$, and therefore
$\dot{m}_{g}(\mbox{${\theta}$}_{o},g_{o})[\widehat{g}-g_{o}]=0.$
@@
$\mathbb{P}\bigl\{r_{\widehat{\theta},\widehat{g}}-r_{\theta_{o},g_{o}}\bigr\}=-\,I(_{o})(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})+o_{p}(n^{-1/2})\,.$
Substituting this into (23) yields
@@
Let $\delta_{n}=\max_{j}(t_{j}-t_{j-1})$ be the mesh size of the time grid. Under Assumptions A1A4, there exist constants $C_{1},C_{2}>0$ such that, for all $n$ and all $\eta$ in the neighborhood $\mathcal{N}_{n}=\{\eta:d(\eta,\eta_{o})\leq M\tau_{n}\}$,
$\sup_{\eta\in\mathcal{N}_{n}}\bigl|\widetilde{\ell}_{n}(\eta)-\ell_{n}(\eta)\bigr|\leq C_{1}\,\delta_{n}$ (24)
@@
$\sup_{\eta\in\mathcal{N}_{n}}\Bigl\|\nabla_{\theta}\widetilde{\ell}_{n}(\eta)-\nabla_{\theta}\ell_{n}(\eta)\Bigr\|\leq C_{2}\,\delta_{n}\,$, (25)
where $\nabla_{\theta}\widetilde{\ell}_{n}(\eta)$ and $\nabla_{\theta}\ell_{n}(\eta)$ are the $\theta$}-score functions of $\widetilde{\ell}_{n}(\eta)$ and $\ell_{n}(\eta)$, respectively.
The proof of Lemma 8 follows directly from classical approximation results for Riemann sums. Under Assumptions A1–A4, $g(t,{\bf x})$ is uniformly Hölder continuous in $t$, uniformly over $\bf x$} and over $\eta\in\mathcal{N}_{n}$. Because the hazard is Lipschitz in $g(t,{\bf x})$, the integrand $Y_{i}(t)\exp\{g(t,{\bf X}_{i})+\mbox{${\theta}$}^{\top}{\bf Z}_{i}\}$ is Lipschitz on each interval where $Y_{i}(t)=1$. Therefore the Riemann approximation error is $O(\delta_{n})$. Applying this bound uniformly over all observations and all $\eta\in\mathcal{N}_{n}$ yields inequalities (24) and (25). Because the argument involves no stochastic elements and is fully deterministic, we omit the proof.
@@
Let $S_{n}(\eta)=\nabla_{\theta}\ell_{n}(\eta)$, and $\dot{S}_{n}(\eta)=\nabla_{\theta}S_{n}(\eta)$ is the Hessian of the empirical log–likelihood with respect to $\theta$}. For a fixed $M>0$, define the shrinking neighbourhood $\mathcal{N}_{n}=\bigl\{\eta:d(\eta,\eta_{o})\leq M\tau_{n}\bigr\}$, $\tau_{n}=\gamma_{n}\log^{2}n$. Assume A1A4. Then,
$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\dot{S}_{n}(\eta)+I(\mbox{${\theta}$}_{o})\bigr\|\;\xrightarrow{P}\;0\,$,
where $I(\mbox{${\theta}$}_{o})$ is the efficient information matrix defined in Theorem 3, and here $\|\cdot\|$ may be any fixed matrix norm; for concreteness, and because it provides the appropriate control on the inverse Hessian, we take it to be the operator norm induced by the Euclidean norm.
@@
$\dot{S}_{n}(\eta)=\mathbb{P}_{n}\dot{s}(\mbox{${\theta}$},g)$
and
$\mathbb{P}\,\dot{s}(\mbox{${\theta}$},g)=E_{0}\!\left[-\int_{0}^{\tau}{\bf Z}{\bf Z}^{\top}Y(t)\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\,dt\right],$
where $\mathbb{P}_{n}$ is the empirical measure and $\mathbb{P}=P_{0}$ denotes expectation under the true law.
Define the matrix-valued per-observation information integrand by
$J(\eta;{\bf W})=\int_{0}^{\tau}{\bf Z}{\bf Z}^{\top}Y(t)\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\,dt,$
where ${\bf W}=(T,\Delta,{\bf X},{\bf Z})$ so that $\dot{S}_{n}(\eta)=-\,\mathbb{P}_{n}J(\eta;\cdot)$ and $\mathbb{P}\dot{s}(\mbox{${\theta}$},g)=-\,\mathbb{P}J(\eta;\cdot)$.

### v2

@@ Appendix A Proofs for Section 4
Let ${\bf u}\in\mathbb{R}^{p}$ and let $\{g_{s}:s\in(-1,1)\}\in\mathcal{H}_{g_{o}}$ be a one-dimensional regular submodel through $g_{o}$, with
$\dot{g}_{0}=\left.\frac{\partial g_{s}}{\partial s}\right|_{s=0}\in\mathcal{T}_{g_{o}}.$
@@
$\{\mathbb{P}_{\theta_{o}+s{\bf u},\;g_{s}}:s\in(-1,1)\}.$
For one observation $(T,\Delta,{\bf X},{\bf Z})$,
@@
Differentiating at $s=0$ gives
$\left.\frac{d}{ds}\ell(\mbox{${\theta}$}_{o}+s{\bf u}\,,\,g_{s})\right|_{s=0}={\bf u}^{\top}\dot{\ell}_{\theta_{o}}+\dot{\ell}_{g_{o},\dot{g}_{0}},$
@@
$\dot{\ell}_{\theta_{o}}=\int_{0}^{\tau}{\bf Z}\,dM(t),\qquad\dot{\ell}_{g_{o},g}=\int_{0}^{\tau}g(t,{\bf X})\,dM(t).$
Hence the nuisance score space is
$\mathcal{T}^{\mathrm{nuis}}_{g_{o}}=\{\dot{\ell}_{g_{o},g}:g\in\mathcal{T}_{g_{o}}\},$
and the efficient score is
@@
where $\Pi_{g_{o}}$ denotes projection onto $\overline{\mathcal{T}^{\mathrm{nuis}}_{g_{o}}}$ in $L_{2}(\mathbb{P}_{0})$.
By Lemma 1 of Sasieni (1992), for any square-integrable predictable functions $\mu_{1},\mu_{2}$,
$E_{0}\!\left[\left(\int_{0}^{\tau}\mu_{1}\,dM\right)\left(\int_{0}^{\tau}\mu_{2}\,dM\right)\right]=E_{0}\!\left[\Delta\,\mu_{1}(T,{\bf X})\mu_{2}(T,{\bf X})\right].$
Therefore the map
$J:h\mapsto\int_{0}^{\tau}\mu(t,{\bf X})\,dM(t)$
is an isometry from $L_{2}(\mathbb{P}_{0},\Delta)$ onto its image. Since $\overline{\mathcal{T}}_{g_{o}}$ is a closed linear subspace of $L_{2}(\mathbb{P}_{0},\Delta)$ and $Z_{j}\in L_{2}(\mathbb{P}_{0},\Delta)$ by Assumption A2, the Hilbert projection theorem implies that for each $j=1,\dots,p$ there exists a unique $g_{j}^{*}\in\overline{\mathcal{T}}_{g_{o}}$ such that
$\Pi_{g_{o}}(\dot{\ell}_{\theta_{o},j})=\dot{\ell}_{g_{o},g_{j}^{*}}.$
Let ${\bf g}^{*}=(g_{1}^{*},\dots,g_{p}^{*})^{\top}$. Then
@@
The orthogonality condition for the projection is
$E_{0}\!\left[\{\dot{\ell}_{\theta_{o},j}-\dot{\ell}_{g_{o},g_{j}^{*}}\}\dot{\ell}_{g_{o},g}\right]=0\qquad\forall g\in\mathcal{T}_{g_{o}},$
which, by the isometry above, is equivalent to
$E_{0}\!\left[\Delta\{Z_{j}-g_{j}^{*}(T,{\bf X})\}g(T,{\bf X})\right]=0\qquad\forall g\in\mathcal{T}_{g_{o}}.$
By density, the same holds for all $g\in\overline{\mathcal{T}}_{g_{o}}$. Thus $g_{j}^{*}$ is the $\Delta$-weighted $L_{2}(\mathbb{P}_{0})$ projection of $Z_{j}$ onto $\overline{\mathcal{T}}_{g_{o}}$, and equivalently
${\bf g}^{*}\in\arg\min_{{\bf g}\in(\overline{\mathcal{T}}_{g_{o}})^{p}}E_{0}\!\left[\Delta\|{\bf Z}-{\bf g}(T,{\bf X})\|_{2}^{2}\right].$
Finally, by the martingale isometry,
$I(\mbox{${\theta}$}_{o})=E_{0}\!\left[\ell_{\theta_{o}}^{*\otimes 2}\right]=E_{0}\!\left[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}^{\otimes 2}\right].$
Finally, under the regularity conditions in Assumptions (A1)–(A4), existence and uniqueness of the weighted projection $\bf g$$^{*}$ in $\overline{\mathcal{T}}_{g_{o}}^{p}$ follow from the approximation theorems of Stone (1985) and the general semiparametric projection theory in Appendix A.4 of Bickel et al. (1993). Therefore, both the efficient score and the efficient information matrix are well-defined and given by the formulas above. ∎
@@
where ${\bf g}^{*}=(g_{1}^{*},\dots,g_{p}^{*})^{\top}$ is the pp–dimensional minimizer from Theorem 3, which depends only on $(\mbox{${\theta}$}_{o},g_{o})$. By Theorem 3, at the true parameter $(\mbox{${\theta}$}_{o},g_{o})$ this coincides with the efficient score,
@@
The empirical efficient score at $(\mbox{${\theta}$},g)$ is $\mathbb{P}_{n}r_{\theta,g}$.
In the following it will be shown that $\mathbb{P}_{n}r_{\widehat{\theta},\widehat{g}}=\mathbf{0}$. For the $j$-th coordinate of $r_{\theta,g}$, we have
$r_{\theta,g}(T,{\bf X},{\bf Z},\Delta;j)$ $=\Delta\{Z_{j}-g_{j}^{*}(T,{\bf X})\}-\int_{0}^{\tau}\{Z_{j}-g_{j}^{*}(t,{\bf X})\}Y(t)\exp\{g(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}\,dt$
$=\left[\Delta Z_{j}-\int_{0}^{\tau}Z_{j}\,Y(t)\exp\{g(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}\,dt\right]$
$\qquad-\left[\Delta g_{j}^{*}(T,{\bf X})-\int_{0}^{\tau}g_{j}^{*}(t,{\bf X})\,Y(t)\exp\{g(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}\,dt\right].$
Therefore,
$\mathbb{P}_{n}\!\bigl[r_{\theta,g}(\cdot\,;j)\bigr]$ $=\mathbb{P}_{n}\!\left[\Delta Z_{j}-\int_{0}^{\tau}Z_{j}\,Y(t)\exp\{g(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}\,dt\right]$
$\qquad-\mathbb{P}_{n}\!\left[\Delta g_{j}^{*}(T,{\bf X})-\int_{0}^{\tau}g_{j}^{*}(t,{\bf X})\,Y(t)\exp\{g(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}\,dt\right]$
$=\dot{\ell}_{n,\theta,j}(\mbox{${\theta}$},g)-\dot{\ell}_{n,g}(\mbox{${\theta}$},g)[g_{j}^{*}],$
where
$\dot{\ell}_{n,\theta}(\mbox{${\theta}$},g)=\mathbb{P}_{n}\!\left[\Delta{\bf Z}-\int_{0}^{\tau}{\bf Z}\,Y(t)\exp\{g(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}\,dt\right],$
and, for any admissible nuisance direction $\mu\in\overline{\mathcal{T}}_{g_{o}}$, the directional Gateaux derivative score
$\dot{\ell}_{n,g}(\mbox{${\theta}$},g)[\mu]=\mathbb{P}_{n}\!\left[\Delta\mu(T,{\bf X})-\int_{0}^{\tau}\mu(t,{\bf X})\,Y(t)\exp\{g(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}\,dt\right].$
[…]
$\mathbb{P}_{n}\,r_{\widehat{\theta},\widehat{g}}(\cdot\,;j)=\dot{\ell}_{n,\theta,j}(\widehat{}\mbox{${\theta}$},\widehat{g})-\dot{\ell}_{n,g}(\widehat{}\mbox{${\theta}$},\widehat{g})[g_{j}^{*}]=0,$
for every $j=1,\dots,p$. Hence $\mathbb{P}_{n}r_{\hat{\theta},\hat{g}}=\mathbf{0}$.
Now, write
@@
$\mathbb{P}\bigl\{r_{\widehat{\theta},\widehat{g}}-r_{\theta_{o},g_{o}}\bigr\}=-\,\mathbb{P}_{n}r_{\theta_{o},g_{o}}+o_{p}(n^{-1/2}).$ (A.14)
@@
Fix a deterministic scalar direction $\mu\in\overline{T}_{g_{o}}$ and consider $g_{s}=g_{o}+s\mu$. Differentiating under the integral sign and applying the martingale compensation identity yields
$\left.\frac{d}{ds}\right|_{s=0}m(\mbox{${\theta}$}_{o},g_{s})=-\,\mathbb{E}\bigl[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}\mu(T,{\bf X})\bigr]\,.$
By Theorem 3, $\bf g$$^{*}$ is the coordinatewise $\Delta$–weighted $L^{2}(\mathbb{P})$–projection of $\bf Z$} onto the nuisance tangent space $(\overline{\mathcal{T}}_{g_{o}})^{p}$. Therefore,
$\mathbb{E}\bigl[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}\mu(T,{\bf X})\bigr]={\bf 0}\quad\text{for all }\mu\in\overline{\mathcal{T}}_{g_{o}}\,$,
and hence $\dot{m}_{g}(\mbox{${\theta}$}_{o},g_{o})[\mu]={\bf 0}$ for every $\mu\in\overline{\mathcal{T}}_{g_{o}}$. Arguing as in Zhong et al. (2022) (Lemma 5), the sieve score equation together with the tangent-space approximation implies that
$\dot{m}_{g}(\mbox{${\theta}$}_{o},g_{o})[\widehat{g}-g_{o}]=o_{p}(n^{-1/2})\,.$
@@
$\mathbb{P}\bigl\{r_{\widehat{\theta},\widehat{g}}-r_{\theta_{o},g_{o}}\bigr\}=-\,I(\theta_{o})(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})+o_{p}(n^{-1/2})\,.$
Substituting this into (A.14) yields
@@
Let $0=t_{0}<t_{1}<\cdots<t_{m_{n}}=\tau$ be a grid with mesh size $\delta_{n}=\max_{1\leq j\leq m_{n}}(t_{j}-t_{j-1})$, and suppose that the grid contains all observed times $T_{1},\dots,T_{n}$. For $\eta=(\mbox{${\theta}$},g)$, let $\mathcal{N}_{n}=\{\eta:d(\eta,\eta_{o})\leq M\tau_{n}\}$ for some fixed $M>0$. Assume A2A3. Also define the uniform time-modulus
$\omega_{n}(\delta)=\sup_{\eta\in\mathcal{N}_{n}}\sup_{{\bf X}\in\mathcal{X}}\sup_{\begin{subarray}{c}s,t\in[0,\tau]\\ |t-s|\leq\delta\end{subarray}}|g(t,{\bf X})-g(s,{\bf X})|.$
Then there exist constants $C_{1},C_{2}>0$, depending only on the bounds in A2A3, such that
$\sup_{\eta\in\mathcal{N}_{n}}\bigl|\tilde{\ell}_{n}(\eta)-\ell_{n}(\eta)\bigr|\leq C_{1}\,\omega_{n}(\delta_{n}),$
@@
$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\nabla_{\theta}\tilde{\ell}_{n}(\eta)-\nabla_{\theta}\ell_{n}(\eta)\bigr\|\leq C_{2}\,\omega_{n}(\delta_{n}).$
In particular, if $\omega_{n}(\delta_{n})=O(\delta_{n})$, then both differences are $O(\delta_{n})$ uniformly on $\mathcal{N}_{n}$.
Proof of Lemma 14.
Fix $\eta=(\mbox{${\theta}$},g)\in\mathcal{N}_{n}$. Recall
$\ell_{n}(\eta)=\frac{1}{n}\sum_{i=1}^{n}\left[\Delta_{i}\chi_{\eta}(T_{i},{\bf X}_{i},{\bf Z}_{i})-\int_{0}^{\tau}Y_{i}(t)e^{\chi_{\eta}(t,{\bf X}_{i},{\bf Z}_{i})}\,dt\right],$
and
$\tilde{\ell}_{n}(\eta)=\frac{1}{n}\sum_{i=1}^{n}\sum_{j:Y_{i}(t_{j-1})=1}\left[\Delta_{ij}\chi_{\eta}(t_{j},{\bf X}_{i},{\bf Z}_{i})-\bigl\{\min(T_{i},t_{j})-t_{j-1}\bigr\}e^{\chi_{\eta}(t_{j},{\bf X}_{i},{\bf Z}_{i})}\right]\,$,
where $\chi_{\eta}(t,{\bf X},{\bf Z})=g(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}$. Because the grid contains all observed times, for each $i$ there exists an index $j(i)$ such that $T_{i}=t_{j(i)}$. Hence
$\sum_{j:Y_{i}(t_{j-1})=1}\Delta_{ij}\chi_{\eta}(t_{j},{\bf X}_{i},{\bf Z}_{i})=\Delta_{i}\chi_{\eta}(T_{i},{\bf X}_{i},{\bf Z}_{i}),$
so the event contribution is represented exactly. Therefore
$\tilde{\ell}_{n}(\eta)-\ell_{n}(\eta)=-\frac{1}{n}\sum_{i=1}^{n}R_{i}(\eta),$
where
$R_{i}(\eta)=\sum_{j:Y_{i}(t_{j-1})=1}\bigl\{\min(T_{i},t_{j})-t_{j-1}\bigr\}e^{\chi_{\eta}(t_{j},{\bf X}_{i},{\bf Z}_{i})}-\int_{0}^{\tau}Y_{i}(t)e^{\chi_{\eta}(t,{\bf X}_{i},{\bf Z}_{i})}\,dt.$
For each $i$, define $\psi_{i,\eta}(t)=Y_{i}(t)e^{\chi_{\eta}(t,{\bf X}_{i},{\bf Z}_{i})}$, $t\in[0,\tau]$. On any interval $(t_{j-1},t_{j}]$ such that $Y_{i}(t_{j-1})=1$, we have $Y_{i}(t)=1$ for $t\in[t_{j-1},\min(T_{i},t_{j})]$, and thus
$\int_{t_{j-1}}^{\min(T_{i},t_{j})}\psi_{i,\eta}(t)\,dt=\int_{t_{j-1}}^{\min(T_{i},t_{j})}e^{\chi_{\eta}(t,{\bf X}_{i},{\bf Z}_{i})}\,dt.$
Since $\Theta$ is compact, $\|{\bf Z}\|\leq B_{{\bf Z}}$ a.s. by A2, and $\|g\|_{\infty}\leq D_{n}=O(1)$ by A3, there exists a constant $C<\infty$ such that uniformly over $i$, $\eta\in\mathcal{N}_{n}$, and $t\in[0,\tau]$,
$e^{\chi_{\eta}(t,{\bf X}_{i},{\bf Z}_{i})}\leq C.$
Moreover, by the mean-value theorem,
$\bigl|e^{\chi_{\eta}(t,{\bf X}_{i},{\bf Z}_{i})}-e^{\chi_{\eta}(s,{\bf X}_{i},{\bf Z}_{i})}\bigr|\leq C\,|g(t,{\bf X}_{i})-g(s,{\bf X}_{i})|\leq C\,\omega_{n}(\delta_{n})$
whenever $|t-s|\leq\delta_{n}$. Therefore, for each interval,
$\left|\int_{t_{j-1}}^{\min(T_{i},t_{j})}e^{\chi_{\eta}(t,{\bf X}_{i},{\bf Z}_{i})}\,dt-\bigl\{\min(T_{i},t_{j})-t_{j-1}\bigr\}e^{\chi_{\eta}(t_{j},{\bf X}_{i},{\bf Z}_{i})}\right|\leq(t_{j}-t_{j-1})\,C\,\omega_{n}(\delta_{n}).$
Summing over $j$ and using $\sum_{j}(t_{j}-t_{j-1})=\tau$, we obtain
$|R_{i}(\eta)|\leq C\tau\,\omega_{n}(\delta_{n})$
uniformly in $i$ and $\eta\in\mathcal{N}_{n}$. Hence
$\sup_{\eta\in\mathcal{N}_{n}}|\tilde{\ell}_{n}(\eta)-\ell_{n}(\eta)|\leq\frac{1}{n}\sum_{i=1}^{n}\sup_{\eta\in\mathcal{N}_{n}}|R_{i}(\eta)|\leq C_{1}\,\omega_{n}(\delta_{n})$
for some constant $C_{1}>0$.
Next consider the $\theta$}-scores. Since $g$ does not depend on $\theta$,
$\nabla_{\theta}\ell_{n}(\eta)=\frac{1}{n}\sum_{i=1}^{n}\left[\Delta_{i}{\bf Z}_{i}-\int_{0}^{\tau}{\bf Z}_{i}Y_{i}(t)e^{\chi_{\eta}(t,{\bf X}_{i},{\bf Z}_{i})}\,dt\right],$
and
$\nabla_{\theta}\tilde{\ell}_{n}(\eta)=\frac{1}{n}\sum_{i=1}^{n}\left[\Delta_{i}{\bf Z}_{i}-\sum_{j:Y_{i}(t_{j-1})=1}\bigl\{\min(T_{i},t_{j})-t_{j-1}\bigr\}{\bf Z}_{i}e^{\chi_{\eta}(t_{j},{\bf X}_{i},{\bf Z}_{i})}\right].$
Therefore,
$\nabla_{\theta}\tilde{\ell}_{n}(\eta)-\nabla_{\theta}\ell_{n}(\eta)=-\frac{1}{n}\sum_{i=1}^{n}{\bf Z}_{i}R_{i}(\eta).$
Using $\|{\bf Z}_{i}\|\leq B_{Z}$ a.s. from A2 together with the bound on $R_{i}(\eta)$,
$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\nabla_{\theta}\tilde{\ell}_{n}(\eta)-\nabla_{\theta}\ell_{n}(\eta)\bigr\|\leq\frac{1}{n}\sum_{i=1}^{n}\|{\bf Z}_{i}\|\,\sup_{\eta\in\mathcal{N}_{n}}|R_{i}(\eta)|\leq B_{{\bf Z}}C_{1}\,\omega_{n}(\delta_{n}).$
This proves the second bound with $C_{2}=B_{{\bf Z}}C_{1}$. ∎
@@
Let $S_{n}(\eta)=\nabla_{\theta}\ell_{n}(\eta)$ and $\dot{S}_{n}(\eta)=\nabla_{\theta}S_{n}(\eta)$ be the Hessian of the empirical log–likelihood with respect to $\theta$}. For a fixed $M>0$, define the shrinking neighbourhood $\mathcal{N}_{n}=\{\eta:d(\eta,\eta_{o})\leq M\tau_{n}\}$, where $\tau_{n}=\gamma_{n}\log^{2}n$. Assume A1A4. Then,
$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\dot{S}_{n}(\eta)+I_{0}(\mbox{${\theta}$}_{o})\bigr\|\;\xrightarrow{P}\;0,$
where
$I_{0}(\mbox{${\theta}$}_{o})=\mathbb{E}_{0}\!\left[\int_{0}^{\tau}{\bf Z}{\bf Z}^{\top}Y(t)\exp\{\mbox{${\theta}$}_{o}^{\top}{\bf Z}+g_{o}(t,{\bf X})\}\,dt\right]=\mathbb{E}_{0}\!\left[\Delta\,{\bf Z}{\bf Z}^{\top}\right].$
If, in addition, $I_{0}(\mbox{${\theta}$}_{o})$ is nonsingular, then
$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\dot{S}_{n}(\eta)^{-1}\bigr\|=O_{p}(1).$
@@
$\dot{S}_{n}(\eta)=\mathbb{P}_{n}\dot{s}(\mbox{${\theta}$},g).$
Define the matrix-valued per-observation integrand
$J(\eta;{\bf W})=\int_{0}^{\tau}{\bf Z}{\bf Z}^{\top}Y(t)\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\,dt,\qquad{\bf W}=(T,\Delta,{\bf X},{\bf Z}),$
so that
$\dot{S}_{n}(\eta)=-\,\mathbb{P}_{n}J(\eta;\cdot),\qquad\mathbb{P}\dot{s}(\mbox{${\theta}$},g)=-\,\mathbb{P}J(\eta;\cdot).$